OpenStudy (heart_offire):

Determine the range of the function. (0,2) (2,4) (4,6) (6, 8) (8, 10) A) y ≤ 10 B) {2, 4, 6, 8, 10} C) 2 ≤ y ≤ 10 D) {0, 2, 4, 6, 8, 10}

OpenStudy (er.mohd.amir):

domain is first elements of ordered pair and range is second elements of ordered pair

OpenStudy (heart_offire):

i know im just confused as to why there are greater and less than signs

OpenStudy (er.mohd.amir):

{2,4,6,8,10}

OpenStudy (er.mohd.amir):

also 2\[2 \le y \le10\]

OpenStudy (er.mohd.amir):

I is written set notation and II written in interval notation

OpenStudy (zzr0ck3r):

\(2\le x \le 10\) has an infinite amount of numbers. I would assume they are working with the reals

OpenStudy (zzr0ck3r):

also, that is called compound inequality not interval notation

Vocaloid (vocaloid):

the ones with ≤ signs can't be your answers, since these answers include numbers that are not whole numbers

Vocaloid (vocaloid):

your range is just the list of y-values from the points in the problem